# Parallel transport on matrix manifolds and Exponential Action

### v1

@@ 6. A quotient metric
Assume the connected subgroup $\mathrm{K}=\mathrm{K}_{\mathfrak{a}}\times\mathrm{K}_{\mathfrak{a}_{{}_{\top}}}\subset\mathrm{G}$ with Lie algera $\mathfrak{k}=\mathfrak{k}_{\mathfrak{a}}\oplus\mathfrak{k}_{\mathfrak{a}_{{}_{\top}}}$ satisfies
(37) $\mathfrak{k}_{\mathfrak{a}}\subset\mathfrak{a};\mathfrak{k}_{\mathfrak{a}_{{}_{\top}}}\subset\mathfrak{a}_{{}_{\top}}\text{ and }\mathfrak{k}^{\mathsf{T}}=\mathfrak{k}$
and $\mathrm{K}$ acts smoothly and properly on $\mathrm{G}$ by right multiplication. Then $\mathrm{K}$ acts as isometries of the metric in theorem 1, and $\mathcal{M}=\mathrm{G}/\mathrm{K}$ can be given a quotient metric. Let $\mathfrak{m}=\mathfrak{k}_{{}_{\perp}}$. For $X\in\mathrm{G}$, the vertical space $\mathcal{V}_{X}$ could be identified with $X\mathfrak{k}$, the horizontal space could be identified with $X\mathfrak{m}$. For two horizontal vector fields $\mathtt{Y},\mathtt{Z}$, the horizontal connection is given by $\nabla^{\mathcal{H}}_{\mathtt{Y}}\mathtt{Z}=\mathrm{D}_{\mathtt{X}}\mathtt{Y}+\Gamma^{\mathcal{H}}(\mathtt{X},\mathtt{Y})$, where the horizontal Christoffel function $\Gamma^{\mathcal{H}}$ for two horizontal vectors $\xi,\eta$ is given by
(38) $\begin{gathered}\Gamma^{\mathcal{H}}(\xi,\eta)=-\frac{1}{2}\left(\xi X^{-1}\eta+\eta X^{-1}\xi+X[X^{-1}\xi,X^{-1}\eta]_{\mathfrak{k}}\right)\\ +\frac{1+\beta}{2}X\left([(X^{-1}\xi)_{\mathfrak{a}},X^{-1}\eta]+[(X^{-1}\eta)_{\mathfrak{a}},X^{-1}\xi]\right).\end{gathered}$
The horizontal geodesic $\gamma$ with $\gamma(0)=X,\dot{\gamma}(0)=\xi$, where $\xi$ is a horizontal vector is given by eq. 20. Let $\mathtt{a}=X^{-1}\xi$. The parallel transport of a horizontal vector $\eta\in T_{X}\mathcal{M}$ is given by
@@
(39) $\begin{gathered}\mathtt{T}(\gamma)_{0}^{t}\eta=X\operatorname{expm}(t\left(-(1+\beta)\mathtt{a}_{\mathfrak{a}}+\mathtt{a}\right))W(t)U(t)\end{gathered}$
where $U(t):=\operatorname{expm}(t(1+\beta)\mathtt{a}_{\mathfrak{a}})$ and $W(t)$ is a solution of the equation $W(0)=X^{-1}\eta$,
@@
(40) $\dot{W}=\frac{1}{2}\left([W,\mathtt{a}]-U(U^{-1}[W,\mathtt{a}]U)_{\mathfrak{k}}U^{-1}+(1+\beta)([\mathtt{a}_{\mathfrak{a}},W]-[W_{\mathfrak{a}},\mathtt{a}])\right).$
Assume further that $(U^{-1}[W,\mathtt{a}]U)_{\mathfrak{k}}=U^{-1}[W,\mathtt{a}]_{\mathfrak{k}}U$, then $W=\operatorname{expv}(\mathrm{P}_{\mathtt{a}},X^{-1}\eta)$ where for $\mathtt{b}\in\mathfrak{g}$
(41) $\mathrm{P}_{\mathtt{a}}:\mathtt{b}\mapsto\frac{1}{2}\left([\mathtt{b},\mathtt{a}]_{\mathfrak{m}}+(1+\beta)([\mathtt{a}_{\mathfrak{a}},\mathtt{b}]-[\mathtt{b}_{\mathfrak{a}},\mathtt{a}])\right).$
This condition is satisfied if $\beta=-1$ or $\mathfrak{k}\subset\mathfrak{a}_{{}_{\top}}$. In these two cases, $\mathrm{P}_{\mathfrak{a}}$ maps $\mathfrak{m}$ to itself and the operator $\mathrm{P}_{\mathfrak{a}}$ is antisymmetric in the pairing $\langle\rangle_{\vec{\beta}}$ on $\mathfrak{m}$.
We note the horizontal Christoffel function differs from the tangent one eq. 18 by a vertical term $\frac{1}{2}X[X^{-1}\xi,X^{-1}\eta]_{\mathfrak{k}}$.
We will see the condition for the simplified parallel transport formula is satisfied for Stiefel manifolds.
@@
As $\mathfrak{a}$ and $\mathfrak{a}_{{}_{\top}}$ commute, $\mathrm{K}_{\mathfrak{a}}$ and $\mathrm{K}_{\mathfrak{a}_{{}_{\top}}}$ also commute, thus, we only need to prove each subgroup acts as isometries. For $\mathtt{k}\in\mathfrak{k}$, set $K=\exp(\mathtt{k})\in\mathrm{K}$. For $\eta=X\mathtt{g}$ in $T_{X}\mathrm{G}$ with $\mathtt{g}\in\mathfrak{g}$, right multiplication by $K$ sends $X$ to $XK$ and $\eta$ to $\eta K$. Hence,
@@
$\begin{gathered}\langle\eta K,\eta K\rangle_{XK,\vec{\beta}}=\langle(XK)^{-1}X\mathtt{g}K,(XK)^{-1}X\mathtt{g}K\rangle_{\vec{\beta}}\\ =\beta_{0}\operatorname{Tr}(K^{-1}\mathtt{g}K)^{2}-\beta_{0}(1+\beta)\operatorname{Tr}((K^{-1}\mathtt{g}K)_{\mathfrak{a}})^{2}.\end{gathered}$
Note $K^{-1}\mathtt{g}K=\operatorname{expv}(-\operatorname{ad}_{\mathtt{k}},\mathtt{g})$. If $\mathtt{k}\in\mathfrak{k}_{\mathfrak{a}_{{}_{\top}}}$ then $\operatorname{ad}_{\mathtt{k}}\mathfrak{g}\subset\mathfrak{a}_{{}_{\perp}}$. Hence, $(\operatorname{ad}_{\mathtt{k}}^{j}\mathtt{g})_{\mathfrak{a}}=0$ for $j>0$, and $(K^{-1}\mathtt{g}K)_{\mathfrak{a}}=\mathtt{g}_{\mathfrak{a}}$. Thus, $\mathrm{K}_{\mathfrak{a}_{{}_{\top}}}$ acts as isometries, as the above reduces to
$\beta_{0}\operatorname{Tr}K^{-1}\mathtt{g}^{2}K-\beta_{0}(1+\beta)\operatorname{Tr}\mathtt{g}_{\mathfrak{a}}^{2}=\beta_{0}\operatorname{Tr}\mathtt{g}^{2}-\beta_{0}(1+\beta)\operatorname{Tr}\mathtt{g}_{\mathfrak{a}}^{2}.$
Similarly, if $\mathtt{k}\in\mathfrak{k}_{\mathfrak{a}}$ then $\operatorname{ad}_{\mathtt{k}}^{j}\mathtt{g}\subset\operatorname{ad}_{\mathtt{k}}^{j}\mathtt{g}_{\mathfrak{a}}+\mathfrak{a}_{{}_{\Join}}$ for $j>0$ by induction, and $(K^{-1}\mathtt{g}K)_{\mathfrak{a}}=K^{-1}\mathtt{g}_{\mathfrak{a}}K$. Then $\langle\eta K,\eta K\rangle_{XK,\vec{\beta}}$ reduces again to $\beta_{0}\operatorname{Tr}\mathtt{g}^{2}-\beta_{0}(1+\beta)\operatorname{Tr}\mathtt{g}_{\mathfrak{a}}^{2}$. Hence, $\mathrm{K}_{\mathfrak{a}}$ and thus $\mathrm{K}$ are groups of isometries.
We need to prove for two horizontal vector fields $\mathtt{Y},\mathtt{Z}$, $\nabla^{\mathcal{H}}_{\mathtt{Y}}\mathtt{Z}$ with $\Gamma^{\mathcal{H}}$ given in eq. 38 is horizontal, metric compatible and projects to a torsion-free connection in the quotient. The latter is true as the only nonsymmetric term $\frac{1}{2}X[X^{-1}\xi,X^{-1}\eta]_{\mathfrak{k}}$ is vertical. The second line in eq. 38 is in $\mathfrak{a}_{{}_{\Join}}$, thus, is horizontal since $\mathfrak{k}\subset\mathfrak{a}\oplus\mathfrak{a}_{{}_{\top}}$. If $\mathtt{Y}(X)$ and $\mathtt{Z}$ are horizontal, with $\mathtt{Y}(X)=X\mathtt{y}(X),\mathtt{Z}(X)=X\mathtt{z}(x)$ for $\mathtt{y}(X),\mathtt{z}(X)\in\mathfrak{m}$ then
$\begin{gathered}\mathrm{D}_{\mathtt{Y}}\mathtt{Z}-\frac{1}{2}\left((\mathtt{Y}X^{-1}\mathtt{Z}+\mathtt{Z}X^{-1}\mathtt{Y})+X[X^{-1}\mathtt{Y},X^{-1}\mathtt{Z}]_{\mathfrak{k}}\right)\\ =X\mathtt{y}\mathtt{z}+X\mathrm{D}_{\mathtt{Y}}\mathtt{z}-\frac{1}{2}\left((X\mathtt{y}\mathtt{z}+X\mathtt{z}\mathtt{y})+X[\mathtt{y},\mathtt{z}]_{\mathfrak{k}}\right)\\ =X\mathrm{D}_{\mathtt{Y}}\mathtt{z}+\frac{1}{2}X[\mathtt{y},\mathtt{z}]-\frac{1}{2}X[\mathtt{y},\mathtt{z}]_{\mathfrak{k}}=X\mathrm{D}_{\mathtt{Y}}\mathtt{z}+\frac{1}{2}X[\mathtt{y},\mathtt{z}]_{\mathfrak{m}}.\end{gathered}$
The last expression is horizontal because $\mathrm{D}_{\mathtt{Y}}\mathtt{z}$ is horizontal by linearity. Finally, metric compatibility with horizontal vector fields follows since the extra term $\frac{1}{2}X[\mathtt{y},\mathtt{z}]_{\mathfrak{k}}$ is vertical, which does not contribute when paired with a horizontal vector.
That the horizontal geodesic is a geodesic with initial horizontal velocity is well-known, or we can check the formula directly, as $\frac{1}{2}\gamma[\gamma^{-1}\dot{\gamma},\gamma^{-1}\dot{\gamma}]_{\mathfrak{k}}$ vanishes.
Let $\Delta(t)$ be the transport of $\eta$ along $\gamma(t)$, set $\Delta=\gamma U^{-1}WU$, or $W(t)=U(t)\gamma(t)^{-1}\Delta(t)U(t)^{-1}$, we derive its equation as in theorem 1

### v2

@@ 7. A naturally reductive family of quotient metrics
Assume $\mathcal{A}\subset\mathcal{G}$ are connected matrix Lie groups with Lie algebras $\mathfrak{a}\subset\mathfrak{g}$, with a bi-invariant symmetric bilinear form $\langle,\rangle$ on $\mathfrak{g}$ and $\mathcal{B}$ is an intertwining operator on $\mathfrak{a}$ as in theorem 1. Let $\mathcal{K}$ be a closed Lie group commuting with $\mathcal{A}$, with a nondegenerate Lie algebra $\mathfrak{k}$ of $\mathcal{K}$ satisfying
[…]
Then $\mathcal{K}$ acts by right multiplication as a group of isometries on $\mathcal{G}$ with the metric eq. 18. The resulting quotient space $\mathcal{M}=\mathcal{G}/\mathcal{K}$ with the quotient metric is naturally reductive when identified with $\tilde{\mathcal{G}}/\tilde{\mathcal{K}}$, where $\tilde{\mathcal{G}}=\mathcal{G}\times\mathcal{A}$ and
(44) $\tilde{\mathcal{K}}=\{(KH,H)\in\tilde{\mathcal{G}}|K\in\mathcal{K},H\in\mathcal{A}\}.$
Specifically, let $\mathfrak{m}$ be the orthogonal complement of $\mathfrak{k}$ in $\langle,\rangle$. Then $\mathfrak{a}\subset\mathfrak{m}$. Set
@@
(45) $\begin{gathered}\tilde{\mathfrak{k}}=\{(\mathtt{K}+\mathtt{A},\mathtt{A})|\mathtt{K}\in\mathfrak{k},\mathtt{A}\in\mathfrak{a}\},\\ \tilde{\mathfrak{m}}_{\bar{\mathfrak{q}}}=\{(\mathtt{M},\hat{\mathcal{B}}\mathcal{P}_{\mathfrak{a}}\mathtt{M})|\mathtt{M}\in\mathfrak{m}\}\text{ with }\hat{\mathcal{B}}=(\mathcal{B}+\beta_{0}\operatorname{I}_{\mathfrak{a}})\mathcal{B}^{-1}.\end{gathered}$
Then $\tilde{\mathfrak{g}}=\tilde{\mathfrak{m}}_{\bar{\mathfrak{q}}}\oplus\tilde{\mathfrak{k}}$, with the reductive decomposition for $(\mathtt{X},\mathtt{A})\in\tilde{\mathfrak{g}}$ is given by
@@
(46) $\begin{gathered}(\mathtt{X},\mathtt{A})=\left(\mathcal{P}_{\mathfrak{m}}\mathtt{X}-(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\mathtt{X}+\beta_{0}^{-1}\mathcal{B}\mathtt{A},-(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}(\mathtt{X}-\mathtt{A})\right)\\ +\left(\mathcal{P}_{\mathfrak{k}}\mathtt{X}+\mathtt{A}+(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}(\mathtt{X}-\mathtt{A}),\mathtt{A}+(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}(\mathtt{X}-\mathtt{A})\right).\end{gathered}$
The metric on $\tilde{\mathcal{G}}/\tilde{\mathcal{K}}$ given by eq. 21 restricted to $\tilde{\mathfrak{m}}_{\bar{\mathfrak{q}}}$ corresponds to eq. 18.
When $\operatorname{I}_{\mathfrak{a}}+\beta^{-1}\mathcal{B}$ is invertible, if $\tilde{\mathcal{G}}$ is equipped with the bi-invariant metric eq. 22 then $\mathcal{M}$ is isometric to $\tilde{\mathcal{G}}/\tilde{\mathcal{K}}$. Let $\pi:\mathcal{G}\mapsto\mathcal{G}/\mathcal{K}$ be the quotient submersion and $\mathfrak{q}:\tilde{\mathcal{G}}\mapsto\mathcal{G}$ be the submersion in theorem 1. Then the map
(47) $\bar{\mathfrak{q}}=\pi\circ\mathfrak{q}:\tilde{\mathcal{G}}\mapsto\mathcal{G}/\mathcal{K}=\mathcal{M},\quad\quad\bar{\mathfrak{q}}(X,A)=\llbracket XA^{-1}\rrbracket\quad\text{ for }(X,A)\in\tilde{\mathcal{G}}$
is a pseudo-Riemannian submersion. Here, $\llbracket XA^{-1}\rrbracket\in\mathcal{M}=\mathcal{G}/\mathcal{K}$ is the equivalence class of $XA^{-1}$ defined by the right multiplication by $\mathcal{K}$.
@@
For $K\in\mathcal{K}$, $\nu\in\mathfrak{g}$ and $\mathtt{A}\in\mathfrak{a}$, differentiating $K^{-1}\operatorname{expm}(t\mathtt{A})K=\operatorname{expm}(t\mathtt{A})$, we get $K^{-1}\mathtt{A}K=\mathtt{A}$. Using $\operatorname{Ad}_{K}$ invariance of $\langle,\rangle$
$\begin{gathered}\langle\mathtt{A},K^{-1}\nu K\rangle=\langle K^{-1}\mathtt{A}K,K^{-1}\nu K\rangle=\langle\mathtt{A},\nu\rangle=\langle\mathtt{A},\mathcal{P}_{\mathfrak{a}}\nu\rangle.\end{gathered}$
Thus, $\mathcal{P}_{\mathfrak{a}}K^{-1}\nu K=\mathcal{P}_{\mathfrak{a}}\nu$. For $\eta=X\nu$ in $T_{X}\mathcal{G}$ with $\nu\in\mathfrak{g}$, right multiplication by $K$ sends $X$ to $XK$ and $\eta$ to $\eta K$. By left invariance of $\langle,\rangle_{\vec{\beta}}$ and $\operatorname{Ad}$-invariance of $\langle,\rangle$, if $\vec{\beta}_{XK}\langle,\rangle$ denotes the pseudo-Riemannian product $\langle,\rangle_{\vec{\beta}}$ at XKXK
@@
$\begin{gathered}\vec{\beta}_{XK}\langle\eta K,\eta K\rangle=\vec{\beta}_{XK}\langle X\nu K,X\nu K\rangle=\langle(XK)^{-1}X\nu K,(XK)^{-1}X\nu K\rangle_{\vec{\beta}}\\ =\beta_{0}\langle K^{-1}\nu K,K^{-1}\nu K\rangle-\beta_{0}\langle(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}(K^{-1}\nu K),\mathcal{P}_{\mathfrak{a}}(K^{-1}\nu K)\rangle\\ =\beta_{0}\langle\nu,\nu\rangle-\beta_{0}\langle(\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B})\mathcal{P}_{\mathfrak{a}}\nu,\mathcal{P}_{\mathfrak{a}}\nu\rangle=\langle\nu,\nu\rangle_{\vec{\beta}}.\end{gathered}$
Thus, $\mathcal{K}$ acts as a group of isometries on $\mathcal{G}$ by right multiplication. Hence, $\pi$ is a submersion from $\mathcal{G}$ if $\mathcal{M}$ is equipped with the quotient metric. Since $YKA^{-1}=YA^{-1}K$ for $Y\in\mathcal{G},K\in\mathcal{K}$ and $A\in\mathcal{A}$, there is a well-defined action by $\tilde{\mathcal{G}}$ on $\mathcal{G}/\mathcal{K}$ by $(X,A):\llbracket Y\rrbracket\mapsto\llbracket XYA^{-1}\rrbracket$, and $\tilde{\mathcal{G}}$ acts as a group of isometries. The inverse image of $\llbracket XA^{-1}\rrbracket$ consists of $(X_{1},A_{1})$ such that $X_{1}A_{1}^{-1}=XA^{-1}K$ for $K\in\mathcal{K}$. Set $H=A^{-1}A_{1}$ then $A_{1}=AH$,
$X_{1}=XA^{-1}KA_{1}=XA^{-1}A_{1}K=XHK=XKH$
since $K$ and $H$ commute. Thus, $(X_{1},A_{1})=(X,A)(KH,H)\in(X,A)\tilde{\mathcal{K}}$. Since $\mathcal{K}$ and $\mathcal{H}$ are closed subgroups, $\tilde{\mathcal{K}}$ is closed, identified with the isotropy group at $\llbracket\operatorname{I}_{\mathcal{G}}\rrbracket$. From [33, Proposition 11.12], we can identify $\mathcal{M}=\mathcal{G}/\mathcal{K}$ with $\tilde{\mathcal{G}}/\tilde{\mathcal{K}}$.
We now show $\tilde{\mathfrak{k}}\cap\tilde{\mathfrak{m}}_{\bar{\mathfrak{q}}}=\{0\}$. In eq. 45, for a point in the intersection, $\mathtt{K}=0,\mathtt{A}=\mathtt{M}$ and $\mathtt{M}=\hat{\mathcal{B}}\mathcal{P}_{\mathfrak{a}}\mathtt{M}$, thus $\mathtt{A}=\mathtt{M}=0$. We confirm that we have the equality in eq. 46, and the two terms in the decomposition are in $\tilde{\mathfrak{m}}$ and $\tilde{\mathfrak{k}}$, respectively, giving us the reductive decomposition, where $\operatorname{Ad}_{\tilde{\mathfrak{k}}}$ invariance will follow from the next argument.
We argue as before that we only need to prove the remaining naturally reductive conditions for nonsingular $\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B}$, which follows from bi-invariance of the metric eq. 22 on $\tilde{\mathcal{G}}$. The singular case follows by continuity as the pairing on $\tilde{\mathfrak{m}}_{\bar{\mathfrak{q}}}$ is well-defined and continuous in the parameters. The horizontal lift is still given by eq. 25, restricted to $\mathtt{Z}\in\mathfrak{m}$. Thus, most requirements have been verified in the proof of proposition 3. It remains to show eq. 46 is an orthogonal decomposition with the metric eq. 22 when $\operatorname{I}_{\mathfrak{a}}+\beta_{0}^{-1}\mathcal{B}$ is nonsingular, which is clear. Therefore, the induced metric on $\mathcal{M}$ is naturally reductive. ∎
Lemma 3.
For an intertwining operator $\mathcal{B}$ on $\mathfrak{a}\subset\mathfrak{g}$ with a bi-invariant form $\langle,\rangle$, for $\mathtt{X}_{1},\mathtt{X}_{2}\in\mathfrak{g}$, assume the Lie algebra $\mathfrak{k}$ satisfies eq. 43. Then we have
(48) $\begin{gathered}\langle\mathfrak{a},[\mathcal{B}\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{1},\mathtt{X}_{2}]+[\mathcal{B}\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{2},\mathtt{X}_{1}]\rangle=\{0\},\\ \langle\mathfrak{k},[\mathcal{B}\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{1},\mathtt{X}_{2}]+[\mathcal{B}\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{2},\mathtt{X}_{1}]\rangle=\{0\}.\end{gathered}$
Proof.
This follows since for $\mathtt{A}\in\mathfrak{a}$,
[…]
We have $\langle\mathfrak{k},[\mathcal{B}\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{1},\mathtt{X}_{2}]+[\mathcal{B}\mathcal{P}_{\mathfrak{a}}\mathtt{X}_{2},\mathtt{X}_{1}]\rangle=\{0\}$ by bi-invariance of $\langle,\rangle$ and $\mathfrak{a},\mathfrak{k}]=0$. ∎
From the decomposition $\mathfrak{g}=\mathfrak{m}\oplus\mathfrak{k}$, the horizontal space at $X\in\mathcal{G}$ is $\mathcal{H}_{X}=X\mathfrak{m}$, while the vertical space is $\mathcal{V}_{X}=X\mathfrak{k}$. We now have a quotient version of theorem 2.
Theorem 4.
With the assumption and the metric of theorem 3, the horizontal Christoffel function $\Gamma^{\mathcal{H}}$ for two horizontal vectors $\xi,\eta$ at $X\in\mathcal{G}$ is given by